OpenStudy (anonymous):

Condense 3ln(x+1)-2lny+ln2

OpenStudy (lgbasallote):

step 1: use power rule on all that can

OpenStudy (anonymous):

Can u give me a final answer, i got \[\ln (2(x+1)^{3})\div Y\]

OpenStudy (mertsj):

\[\ln (x+1)^3+\ln 2-\ln y^2=\ln \frac{2(x+1)^3}{y^2}\]